# Misaligned incentives in sports tournaments: A mathematical analysis of the post-2024 UEFA Champions League

### v7

@@ Abstract
UEFA declares that it is fully committed to respecting the fundamental values of sport. However, the rules of the post-2024 UEFA Champions League are shown to be unfair: a match in the 2023/24 German Bundesliga almost created powerful incentives for a team to lose. We develop a mathematical model to reveal how fairness can be ensured by giving priority to domestic leagues in filling vacancies. The implementation of our recommendations reduces or even eliminates any similar problem with misaligned incentives. It is also presented that UEFA repeatedly commits the same mistake by allocating the vacant slots unfairly. Therefore, governing bodies in sports are called to work more closely together with the scientific community.
@@ 1 Introduction
Ensuring the fairness and integrity of sports is a crucial responsibility of tournament organisers (Forrest and McHale,, 2019; Kendall and Lenten,, 2017; Devriesere et al.,, 2024). One of the most obvious criteria of fairness is incentive compatibility, in other words, the lack of misaligned incentives: it should never happen that one agent (player/team) seeks to win and the other to lose. This would be clearly against the spirit of the game and may easily lead to collusion or match-fixing (Csató,, 2024; Haigh,, 2009).
@@
The current paper uncovers that, while UEFA may be committed to sporting merit in principles, their regulations do not always exclude an unfair game with misaligned incentives. Even though the recent reform of the UEFA Champions League improved the seeding system, the assignment of the novel European performance spots may create an incentive to lose. This has almost occurred in the last round of the 2023/24 German Bundesliga in the match Eintracht Frankfurt vs. RB Leipzig. To analyse the issue in depth, we develop a mathematical model to propose two solutions that can improve fairness.
The novel case of incentive incompatibility is especially worrying because “the objective for a sport organisation should be to eliminate any game that raises potential unfairness ex-ante(Scelles et al.,, 2024, p. 535). Furthermore, the root of the problem resides in the allocation of vacant slots, which led to some instances of incentive incompatibility in the past. It seems that filling vacancies poses a significant challenge for tournament organisers, even though several academic papers have been published on this issue (see Section 2 and Table 2). Therefore, governing bodies in major sports, especially UEFA, are called to strengthen collaboration with the scientific community.
@@ 7 SC Freiburg 11 9 13 44 56 --12 42
Table 1 shows the standing of the 2023/24 German Bundesliga on the morning of 18 May 2024 (the last matchday) with a slight modification: the result of the match Borussia Dortmund vs. 1899 Hoffenheim, played on 15 February 2024, is changed from 2-3 to 4-3. At that moment, it was already known that Germany and Italy finished in the top two based on the 2023/24 season association club coefficients, and earned the two European performance spots. The two clubs playing in the final of the UEFA Champions League on 1 June 2024 were known to be Borussia Dortmund and the Spanish club Real Madrid, too.
@@ 4 A mathematical model
The first condition becomes relevant if a club (team $r$) that hopes to have the highest individual club coefficient in the champions path in the next season supports a club (team $j$) from a higher-ranked national association to win the UEFA Champions League in order to create a vacancy to be filled by itself. This is much less threatening than the third condition because tanking is excluded in the final knockout stage and the titleholder usually qualifies via its domestic championship, too. For example, the Ukrainian Shakhtar Donetsk (that had the highest coefficient in the champions path as we have seen in Section 3) played in a group with Antwerp, Barcelona, and Porto in the the 2023/24 Champions League. Therefore, it would be interested in supporting Barcelona to win the Champions League at the expense of Antwerp since it was more likely that Barcelona would also qualify through its domestic championship.
@@ 5 Discussion
  • Rule change A: Modify the order of applying Articles 3.04 and 3.07 in the UEFA Champions League regulation (UEFA, 2024b, ). That is, first the two European performance spots are allocated (UEFA, 2024b, , Article 3.07), and then the possible vacancies created by the titleholders are addressed (UEFA, 2024b, , Article 3.04).
@@ 6 Conclusions and managerial implications
This paper has revealed that the UEFA Champions League rules are incentive incompatible again from the 2024/25 season. In particular, the allocation rules of the vacant slots caused by the titleholders and of the two European performance spots have almost led to a game with misaligned incentives in the 2023/24 German Bundesliga. We have used a formal mathematical model to uncover the underlying causes and to provide two alternative policies: (1) Reverse the order in which the vacancies caused by the titleholders are filled and the European performance sports are allocated; or (2) Fill any vacancy caused by a titleholder within its domestic league. While the first solution substantially reduces the probability of an incentive incompatible situation, the second fully eliminates the problem if the current draw restrictions are retained. Therefore, UEFA is strongly encouraged to adopt one of our recommendations in order to safeguard the integrity of its flagship tournament.
@@
The problem identified is especially worrying because it is rooted in the allocation of vacant slots, a topic that was repeatedly responsible for incentive incompatibility in UEFA club competitions as summarised in Table 2. This issue offers an instructive lesson for decision-makers around the world: commitment to sporting merit requires not only media releases but appropriately designed tournaments. Since rule changes such as the introduction of the European performance spot can always affect fairness, a thorough assessment is necessary before they are implemented.

### v8

# Misaligned incentives in sports: A mathematical analysis of the post-2024 UEFA Champions League qualification

@@ Abstract
UEFA declares that it is committed to respecting the fundamental values of sports. However, the qualification rules of the post-2024 UEFA Champions League are shown to be unfair: a game with misaligned incentives was narrowly avoided in the 2023/24 German Bundesliga. We develop a mathematical model to reveal how incentives for losing can be reduced or eliminated. Since UEFA repeatedly commits the same theoretical mistake in designing the qualification system of its competitions, governing bodies in sports are called to work more closely together with the scientific community.
@@ 1 Introduction
Ensuring the fairness and integrity of sports is a crucial responsibility of tournament organisers (Forrest and McHale,, 2019; Kendall and Lenten,, 2017; Devriesere et al.,, 2024). One of the most important criteria of fairness is incentive compatibility: it should never be allowed to happen that one player or team seeks to win and its opponent to lose. This would be clearly against the spirit of the game and may easily lead to collusion or match-fixing (Csató,, 2024; Haigh,, 2009).
@@
The current paper uncovers that the assignment of the European performance spots, introduced from the 2024/25 season of the UEFA Champions League, may create a misaligned incentive to lose. This was narrowly avoided in the last round of the 2023/24 German Bundesliga, which determined qualification for the next season of the UEFA Champions League. To analyse the issue in depth, we develop a mathematical model to propose two solutions that can improve fairness.
The novel case of incentive incompatibility is especially worrying because “the objective for a sport organisation should be to eliminate any game that raises potential unfairness ex-ante(Scelles et al.,, 2024, p. 535). Furthermore, the same theoretical problem led to some instances of incentive incompatibility in the past, even though several academic papers have been published on this issue (see Section 2 and Table 2). In order to avoid these paradoxes, governing bodies in major sports, especially UEFA, are called to strengthen collaboration with the scientific community.
@@ 3 A real-world example
Since the clubs qualify for the UEFA Champions League based on the results of the previous season, the allocation of the European performance spot could have lead to misaligned incentives in the 2023/24 season of domestic leagues.
@@ 7 SC Freiburg 11 9 13 44 56 --12 42
Table 1 shows the standing of the 2023/24 German Bundesliga on the morning of 18 May 2024 (the last matchday) with a slight modification. In particular, the result of the match Borussia Dortmund vs. 1899 Hoffenheim, played on 15 February 2024, is changed from 2-3 to 4-3. At that moment, it was already known that Germany and Italy finished in the top two based on association club coefficients in the 2023/24 season, and earned the two European performance spots in the 2024/25 Champions League. The two clubs playing in the final of the UEFA Champions League on 1 June 2024 were known to be Borussia Dortmund and the Spanish club Real Madrid, too.
@@ 4 A mathematical model
The first condition becomes relevant if a club (team $r$) that hopes to have the highest individual club coefficient in the champions path in the next season supports a club (team $j$) from a higher-ranked national association to win the UEFA Champions League in order to create a vacancy to be filled by itself. This is much less threatening than the third condition because tanking is excluded in the final knockout stage and the titleholder usually qualifies via its domestic championship, too. For example, the Ukrainian Shakhtar Donetsk (that had the highest coefficient in the champions path as we have seen in Section 3) played in a group with Antwerp, Barcelona, and Porto in the 2023/24 UEFA Champions League. Therefore, it would be interested in supporting Barcelona to win the Champions League at the expense of Antwerp since it was more likely that Barcelona would also qualify through its domestic championship.
@@ 5 Discussion
  • Rule change A: Reverse the order of applying Articles 3.04 and 3.07 in the UEFA Champions League regulation (UEFA, 2024b, ). That is, first the two European performance spots are allocated (UEFA, 2024b, , Article 3.07), and then the possible vacancies created by the titleholders are addressed (UEFA, 2024b, , Article 3.04).
@@ 6 Conclusions and managerial implications
This paper has revealed that the qualification system of the UEFA Champions League is incentive incompatible from the 2024/25 season. In particular, a game with misaligned incentives was narrowly avoided in the 2023/24 German Bundesliga. We have used a formal mathematical model to uncover the underlying causes and to provide two alternative policies. UEFA is strongly encouraged to adopt one of our recommendations in order to safeguard the integrity of its flagship tournament.
@@
The problem identified is especially worrying because it is rooted in an issue (the allocation of vacant slots) that was repeatedly responsible for incentive incompatibility in UEFA club competitions as summarised in Table 2. This offers an instructive lesson for decision-makers around the world: commitment to sporting merit requires not only media releases but also appropriately designed tournaments. Since rule changes can always affect fairness, a thorough assessment is necessary before they are implemented.